Solve the following initial value problems. When it is possible, express the solutions explicitly in terms of the independent variable.

d) *– 4y = 3et, y(0) = -2, y(0) = 0
Answer #1

E Salt andred given d y. -4° = Bet >> (12-4y = Betl You) = -2 3140 =0 put = D- y = cft P.I for cof 024 =0 . D = 12 Cof= get t© Yvo) =0 = 24-262 + from egn ♡ and esh a = – 19716 C = -13/16 ③ y=-*-e*+2tetaplease upvote ????